# a learning based approach to task generalized system identification

### v1

@@ Abstract
We address the problem of robot guided assembly tasks, by using a learning-based approach to identify contact model parameters for known and novel parts. First, a Variational Autoencoder (VAE) is used to extract geometric features of assembly parts. Then, we combine the extracted features with physical knowledge to derive the parameters of a contact model using our newly proposed neural network structure. The measured force from real experiments is used to supervise the predicted forces, thus avoiding the need for ground truth model parameters. Although trained only on a small set of assembly parts, good contact model estimation for unknown objects were achieved. Our main contribution is the network structure that allows us to estimate contact models of assembly tasks depending on the geometry of the part to be joined. Where current system identification processes have to record new data for a new assembly process, our method only requires the 3D model of the assembly part. We evaluate our method by estimating contact models for robot-guided assembly tasks of pin connectors as well as electronic plugs and compare the results with real experiments.
@@ I Introduction
Robots are increasingly being employed for the automated assembly and disassembly of consumer products such as electronic devices as well as industrial components. During an assembly process, two or more components come into contact, see Fig. 1. This results in reaction forces during the joining process. To ensure a reliable assembly process, force-torque controllers such as admittance [1], impedance [2] or hybrid force-position controllers [3] are usually employed. The controller parameters must be designed depending on the contact dynamics of the respective assembly step, therefore needing a suitable contact model. However, a generalization of the model and thus its transferability to similar processes is usually not possible, as only data from a single task is utilized in the design. This implies that the model can only be generalized to unknown input signals, but not to changes in the process information.
@@ II Related Work
Conventional system identification methods capture the behavior of a system by observing its inputs and corresponding outputs. After selecting a suitable model, the input and output data are used in an optimization procedure to determine the model parameters that best fit the system dynamics [4].
@@
Purely neural network based approaches, such as using MLPs [7], CNNs [8] and Autoencoders [9] for model identification can be seen as black box approaches where no prior knowledge is needed. Like in the classical approach, they are optimized on a set of input-output data to minimize the error between real and estimated model response.
@@ III Method
The core idea is to determine the parameters of a model indirectly using the process information $P$, rather than directly with input and output data. Since $P$ can be high dimensional data, it is not directly transferable into a parametric model representation. To solve this issue, we use a neural network to estimate model parameters given current process information $NN(P)\rightarrow\underline{\Theta}$. Where $\dim(\underline{\Theta})$ depends on the model structure $M(\cdot)$ representing the assembly task. This yields a method which is capable of generalizing and thus identifying contact model parameters depending on process information within the related task domain. By change of process information, conventional methods require new measurements to fit a model to these new conditions. In our approach, the model can generalize within the trained task domain, avoiding additional data acquisition. An overview of the general method is given in Fig. 3.
@@ III-A Task as a System
where $x$ and $u$ can be multivariate when used in a multiple-input multiple-output setting. $P=\{\rho_{1},...,\rho_{n}\}$ is a collection of process information influencing the state variable and $M(\underline{\Theta},u)$ can be any appropriate model, here described as a linear differential equation:
@@ IV-C2 Regression Head
The regression head uses the latent-space representation $z$ as input to predict the parameters $\underline{\Theta}$ of the dynamic model. This is done by estimating local models at centers $b$ which are spaced over the value range of $p_{z}$. These $N$ local models are each represented by a Gaussian $\Psi_{n}$ with variance $c$ and amplitude $a$, see Fig. 7. We use softplus as activation for estimating the amplitude and variance to ensure positive non-zero value.
@@
$\underline{\delta}$ is a $N$ dimensional column vector, with $N$ the number of local models and $\underline{\delta_{s}}$ are the scaled sections. $\underline{\delta_{s}}$ are non-overlapping sections in the vertical axis with their lengths correlating with the width of the Gaussian representing the local model. The centers $b$ are then set as the midpoints of $\underline{\delta_{s}}$. This encourages a better distribution of local models over $p_{z}$.
We have to restrict the estimated variances, so that the influence of the local models does not extend over the entire value range of $p_{z}$. This is achieved using (9) to calculate $c$, so that $\Psi(b+\underline{\delta_{s}}/2)=\xi$ applies. $\xi$ is the value of the local model at the edges of $\underline{\delta_{s}}$. This gives us the possibility to constrain the area of influence of the local models. We use $\xi=a/20$ for $c_{min}$ and $\xi=a/3$ for $c_{max}$.
@@ IV-C3 Dynamic Model
For the dynamic model we choose a 2nd order ODE with varying parameters $\underline{\Theta}$, which is excited by $p_{z}$, see (5). The choice of model depends on the complexity of the state variable. If prior knowledge of the system is available, then a more appropriate model can be selected. Since we are using a superposition of local models, we can adapt the model complexity by increasing the number of local models $N$. Solving this equation yields the estimated force $\hat{F}$ during assembly.
@@ IV-D Training
where $X$ and $Y$ are the two point clouds. Afterwards, we freeze the encoder and append the regression head and the dynamic model. The regression head is then trained for another $100$ epochs with batchsize $10$ and learning rate $2\cdot 10^{-4}$. As a loss function we use the mse between measured $F$ and estimated force $\hat{F}$ during assembly:
$L_{reg}=\frac{1}{n}\sum_{i=1}^{n}(F_{i}-\hat{F}_{i})^{2}$ (11)
where $n$ is the number of measurements during execution. Here too we apply Adam for optimization. For each task domain, we use the training data listed in Table I.
@@ V Conclusion & Future Work
In this work, we have demonstrated the feasibility of estimating contact model parameters for robot guided assembly tasks based on the geometry of the components to be assembled. This was achieved by deriving the model from learned contact behavior of assembly tasks using our proposed network structure. Our approach integrates a conventional dynamic model, which provides insights into the learned contact behavior. This is not possible with black box approaches. Additionally, the integration of prior knowledge is possible by appropriately setting the centers of the local Gaussian or by choosing an overall suitable dynamic model $M(\cdot)$. By embedding a contact model in the network structure, and deriving its parameters based on the geometry of the object, our method combines the dynamic informed property of PINNs with the generalization ability of statistical approaches. Although our method delivers good results, it also has its limitations. Since currently only the geometry of the component to be assembled is considered, imprecise models are estimated for components of the same geometry consisting of different materials. In order to address these limitations, future work will explore the combination of additional process information in latent space to account for material properties in addition to geometry. Furthermore, we will use the estimated dynamic model of the assembly task to perform an automatic controller design and transfer the method to other use-cases in order to test the applicability.

### v3

@@ Abstract
We address the problem of robot guided assembly tasks, by using a learning-based approach to identify contact model parameters for known and novel parts. First, a Variational Autoencoder is used to extract geometric features of assembly parts. In addition, we propose a new type of network structure, combining this VAE with physical knowledge through a contact model, allowing the parameters of the model to be derived by the geometric features. The measured force from real experiments is used to supervise the predicted forces, thus avoiding the need for ground truth model parameters. Although trained only on a small set of assembly parts, good contact model estimation for unknown objects were achieved. Our main contribution is a novel network structure which allows to estimate contact models of assembly tasks depending on the geometry of the part to be joined. Where current system identification processes have to record new data for a new assembly process, our method only requires the 3D model of the assembly part. We evaluate our method by estimating contact models for robot-guided assembly tasks of pin connectors as well as electronic plugs and compare the results with real experiments.
@@ I Introduction
Robots are increasingly being employed for the automated assembly and disassembly of consumer products such as electronic devices as well as industrial components. During an assembly process, two or more components come into contact, see Fig. 1. This results in reaction forces during the joining process. To ensure a reliable assembly process, force-torque controllers such as admittance [1], impedance [2] or hybrid force-position controllers [3] are usually employed. In order to achieve a reliable assembly process, the controller parameters must be designed depending on the contact dynamics of the respective assembly step. A model of contact dynamics is necessary to determine control parameters suitably. However, a generalization of the model and thus its transferability to similar processes is usually not possible, as only data from a single task is utilized in the design. This implies that the model can only be generalized to unknown input signals, but not to changes in the process information.
@@ II Related Work
Conventional system identification methods capture the behavior of a system by observing its inputs and corresponding outputs. After selecting a suitable model, the input and output data are used in an optimization procedure to determine the model parameters that best fits the system dynamics [4].
@@
Purely neural network based approaches, such as using MLPs [7], CNNs [8] and Autoencoders [9] for model identification can be seen as black box approaches where no prior knowledge is needed. Like in the classical approach, they are optimized on a set of input output data to minimize the error between real and estimated model response.
@@ III Method
The basic idea is to determine the parameters of a model indirectly using the process information $P$, rather than directly with input and output data. Since $P$ can be high dimensional data, it is not directly transferable into a parametric model representation. To solve this issue, we use a neural network to estimate model parameters given current process information $NN(P)\rightarrow\underline{\Theta}$. Where $\dim(\underline{\Theta})$ depends on the model structure $M(\cdot)$ representing the assembly task. This yields a method which is capable of generalizing and thus identifying contact model parameters depending on process information within the related task domain. By change of process information, conventional methods require new measurements to fit a model to these new conditions. In our approach, the model can generalize within the trained task domain, avoiding renewed data acquisition. An overview of the general method is given in Fig. 3.
@@ III-A Task as a System
where $x$ and $u$ can be multivariate when used in a MIMO setting. $P=\{\rho_{1},...,\rho_{n}\}$ is a collection of process information influencing the state variable and $M(\underline{\Theta},u)$ can be any appropriate model, here described as a linear differential equation:
@@ IV-C2 Regression Head
The regression head uses the latent-space representation $z$ as input to predict the variable parameters $\underline{\Theta}$ of the dynamic model. This is done by estimating local models at centers $b$ which are spaced over the value range of $p_{z}$. These local models are represented by $N$ Gaussian $\Psi_{n}$ with variance $c$ and amplitude $a$, see Fig. 7. We use softplus as activation for estimating the amplitude and variance to ensure positive non-zero value.
@@
$\underline{\delta}$ is a $N$ dimensional column vector, with $N$ the number of local models and $\underline{\delta_{s}}$ are the scaled sections. Multiplication with $Q$ results in centers $b$ of local models constrained to the value range of $p_{z}$:
$b$ $=\underline{\delta_{s}}^{T}\cdot Q$ (8)
$Q_{N,N}$ $=U_{N,N}-\frac{1}{2}I_{N,N}$ (9)
$U$ is an upper triangular matrix of dimension $N\times N$ consisting of ones and $I$ is the identity of dimension $N\times N$. Thus, $Q$ is an upper triangular matrix consisting of ones with the main diagonal set to $\frac{1}{2}$. This encourages a better distribution of local models over $p_{z}$.
We have to restrict the estimated variances, so that the influence of the local models does not extend over the entire value range of $p_{z}$. This is achieved using (11) to calculate $c$, so that $\Psi(\underline{\delta_{s}}^{T}U)=\xi$ applies. $\xi$ is the value of the local model at the edges of $\underline{\delta_{s}}$. This gives us the possibility to constrain the area of influence of the local models. We use $\xi=a/20$ for $c_{min}$ and $\xi=a/3$ for $c_{max}$.
@@ IV-C3 Dynamic Model
For the dynamic model we choose a 2nd order ODE with varying parameters $\underline{\Theta}$, which is excited by $p_{z}$, see (5). The choice of model depends on the complexity of the state variable. If prior knowledge of the system is available, then a more appropriate model can be selected. Because we are using a superposition of local models, we can adapt the model complexity by increasing the number of local models $N$. Solving this equation yields the estimated force $\hat{F}$ during assembly.
@@ IV-D Training
where $X$ and $Y$ are the two point clouds, while $x$ and $y$ represent each point in their respective point cloud. Afterwards, we freeze the encoder and append the regression head and the dynamic model. The regression head is then trained for another $100$ epochs with batchsize $10$ and learning rate $2\cdot 10^{-4}$. As a loss function we use the MSE between measured $F$ and estimated force $\hat{F}$ during assembly
$L_{reg}=\frac{1}{n}e^{T}e$
where $e=(F-\hat{F})$ and $e\in\mathbb{R}^{n}$. Here too we apply Adam for optimization. For each task domain, we use the training data listed in Table I.
@@ V Conclusion & Future Work
In this work, we have demonstrated the feasibility of estimating contact model parameters for robot guided assembly tasks based on the geometry of the components to be assembled. This was achieved by deriving the model from learned contact behavior of assembly tasks using a novel network structure. Our approach integrates a conventional dynamic model, which provides insights into the learned contact behavior. This is not possible with black box approaches. Additionally, the integration of prior knowledge is possible by appropriately setting the centers of the local Gaussian or by choosing an overall suitable dynamic model $M(\cdot)$. By embedding a contact model in the network structure, and deriving its parameters based on the geometry of the object, our method combines the dynamic informed property of PINNs with the generalization ability of statistical approaches. Although our method delivers good results, it also has its limitations. Since currently only the geometry of the component to be assembled is considered, imprecise models are estimated for components of the same geometry consisting of different materials. In order to address these limitations, future work will explore the combination of additional process information in latent space to account for material properties in addition to geometry. Furthermore, we will use the estimated dynamic model of the assembly task to perform an automatic controller design and transfer the method to other use-cases in order to test the applicability.